Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probablility

Expert replies
by Jinglander » Sat Jul 31, 2010 3:21 pm
Can someone tell me where i went wrong here

A club has 10 members. They will first pick a president then a secretary then a treasure. What is the probability that a given person is selected as secretary or treasurer.

I thought about it this way. First you much not be selected as president this is a 9/10 chance. Then you have a 1/9 and 1/8 chance to be selected as one of the other positions. so 9/10*(1/9+ 1/8)
Join the discussion
Source: — Problem Solving |

by KrazyKarl » Sat Jul 31, 2010 4:03 pm
I think that you're missing the fact that there's only a 8/9 chance that you're not selected as the first of the other positions (say, secretary), so the odds are really:

9/10 chance not president
1/9 chance secretary
8/9 chance not secretary * 1/8 chance that, then, you become treasurer

So a 1/9 * 9/10 chance of being secretary, or 1/10

Then a 9/10 * 8/9 * 1/8 chance that you become treasurer

That would be 1/10 + 1/10 = 2/10. Is that the right answer? It almost seems too easy when you look at it like that, but I think that was your mistake before, so this at least accounts for that.
Join the discussion

by Jinglander » Sat Jul 31, 2010 4:10 pm
Thats right but how do you know what order to apply the fractions any order works?
Join the discussion

by KrazyKarl » Sat Jul 31, 2010 4:19 pm
Well if you're multiplying the order shouldn't matter (is that the transitive property? I know there's a name for it from school), so as long as you're setting it up right you should be okay. Glad that was right! I thought I might have missed something but at least thought that I could see where the first mistake was.
Join the discussion

by Ian Stewart » Sat Jul 31, 2010 9:39 pm
Jinglander wrote:Can someone tell me where i went wrong here

A club has 10 members. They will first pick a president then a secretary then a treasure. What is the probability that a given person is selected as secretary or treasurer.

I thought about it this way. First you much not be selected as president this is a 9/10 chance. Then you have a 1/9 and 1/8 chance to be selected as one of the other positions. so 9/10*(1/9+ 1/8)
You can also look at this in a different way: everyone is equally likely to be chosen secretary. So the given person has a 1/10 probability of being selected as secretary. Similarly, everyone is equally likely to be chosen treasurer, so our person has a 1/10 probability of being selected treasurer. So the person has a 1/10 + 1/10 = 2/10 probability of being selected secretary or treasurer.

And KrazyKarl - I think the phrase you were looking for is 'commutative property' :) (that's not a phrase you need to know for the GMAT though)
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion